OpenStudy (anonymous):

In triangle ABC, A=32 degrees, a=8, and b=7. Find C. A 120.4° B 110.7° C 30.4° D 20.7° (it think its 110.7 but I'm not sure)

OpenStudy (anonymous):

is there a figure for this question?

OpenStudy (anonymous):

sin 32/8 = sin b/7 -> Sin B= 7sin32/8

OpenStudy (aum):

Use the law of sines to first figure out angle B. a / sin(A) = b / sin(B) sin(B) = b / a * sin(A) Plug in the numbers and find B Then C = 180 - (A + B)

OpenStudy (anonymous):

would A be 32?

OpenStudy (aum):

What do you get for angle B? sin(B) = 7/8 * sin(32) = ? Then take the sine inverse to find B.

OpenStudy (anonymous):

inverse meaning 8/7?

OpenStudy (anonymous):

0.46?

OpenStudy (aum):

sin(B) = 7/8 * sin(32) = 0.4637 (this is the sine value) To find angle B you will have to find \(\Large \sin^{-1}(0.4637)\)

OpenStudy (aum):

That is called the sine inverse.

OpenStudy (anonymous):

soooo 36+0.4637-180 = -143?

OpenStudy (aum):

No. That 0.4637 is the sine value. That is NOT the angle B. It is sin(B). sin(B) = 0.4637 Find angle B. B = inverse sine of (0.4637) In your calculator find how to calculate the inverse sine.

OpenStudy (anonymous):

27.6?

OpenStudy (aum):

Yes. Angle B = 27.6 degrees. Angle C = 180 - Angle A - Angle B = 180 - 32 - 27.6 = ?

OpenStudy (anonymous):

120.4? '